Can you please help me with this questions?

1). Use point-slope form to write an equation of the line. Write the equation in standard form. The line passes through (6, −5) and (4, 1).

2). Let f(x) = 7x and g(x) = x2 + 4. Find the function. g ∘ f
Answer #1

(X, Y) = (4,1) 1) Given points rel (x, y) = 6,-5) The ear of line passing the the through (6,5) G (4,1) hay slope Y Y m- 2 Xy = - 3x + 13 2 f(x) = 7% g(x)=x²+4 Now g of (x) = g(f(x)) - 9(7x) = (7x) 44 g of (x) = 498 +4